# Radar-Informed 3D Multi-Object Tracking under Adverse Conditions

### 15 Apr 2026 v1

@@ Abstract
The challenge of 3D multi-object tracking (3D MOT) is achieving robustness in real-world applications, for example under adverse conditions and maintaining consistency as distance increases. To overcome these challenges, sensor fusion approaches that combine LiDAR, cameras, and radar have emerged. However, existing multi-modal fusion methods usually treat radar as another learned feature inside the network. When the overall model degrades in difficult environmental conditions, the robustness advantages that radar could provide are also reduced. In this paper we propose RadarMOT, a radar-informed 3D MOT framework that explicitly uses radar point cloud data as additional observation to refine state estimation and recover detector misses at long ranges. Evaluations on the MAN-TruckScenes dataset (TruckScenes) show that RadarMOT consistently improves the Average Multi-Object Tracking Accuracy (AMOTA) with absolute 12.7% at long range and 10.3% in adverse weather. The code will be available at https://github.com/bingxue-xu/radarmot
@@ I Introduction
Safe 3D perception requires a system to continuously understand what is around it and where they are going before acting. This is the task of 3D 3D MOT: maintaining identities and motion trajectories of detected objects across time. Trajectory prediction and motion planning depend directly on the quality of these estimates. 3D MOT operates by detecting objects independently in each frame and then associating them across time to form trajectories. This makes 3D MOT performance fundamentally dependent on two things: (i) how accurately their motion is estimated (state estimation accuracy), and (ii) how reliably detections are matched with tracked objects (data association robustness).
@@
To reduce the computation burden on the graphics processing unit (GPU), we revisit both the radar sensing mechanism and Kalman filter-based 3D MOT frameworks [33, 31, 3, 24, 43]. Ther, we first incorporate radar point clouds as an additional measurement in the Kalman filter update, which consistently boosts AMOTA and lowers false positives, indicating a more accurate state estimation. Next, we address motion blur during data preprocessing by leveraging radial velocity. Based on the observation that the distance-based tracker CenterPoint [37] yields more true positives than MCTrack [31], we further extend data association into a two-stage scheme that utilizes bidirectional distances with radar association, thereby stabilizing the tracking system.
RadarMOT achieves strong and consistent improvements on TruckScenes, demonstrating that leveraging radar physical measurements in the tracking stage can substantially increase robustness when LiDAR and cameras degrade. Overall, RadarMOT improves glsamota by +6.7% over the MCTrack baseline and reduces glsids by 30%, with the largest gains appearing at long range and under adverse conditions. The contributions of this paper are:
@@
  • We formulate a radar-informed Kalman filter update that uses associated radar radial velocities to refine tracked objects’ planar velocity and stabilize trajectories.
@@ II RELATED WORK
Tracking by Detection TBD divides the tracking task into four independently modules: detection, motion model, data association, and lifecycle management [24]. Since AB3DMOT [33] established a strong baseline by coupling Kalman filter [25] with Intersection over Union (IoU)-based Hungarian matching [9]. The following methods have progressively refined the association strategy, from generalized IoU to normalization [47], lifting it to 3D [24], and combining distance with  IoU [31]. Another representative approach, CenterPoint [37], replaces  IoU with Euclidean distance combined with greedy nearest-point matching, achieving strong performance on nuScenes[2] and Waymo [27]. The following work [3] replaces Euclidean distance by Mahalanobis distance. Other methods [13, 12, 43, 16] enhanced the motion model within the Kalman filter. SimpleTrack [24] identifies non-maximum suppression (NMS) and proposes two-stage association. MCTrack [31], consolidates these advances into a unified and general framework that refines the Kalman filter and incorporates both  IoU and distance into the data association stage, achieving state-of-the-art performance across the KITTI [6], nuScenes [2], and Waymo [27] datasets. We therefore adopt MCTrack as our baseline.
@@ III-A Motion Compensation
Ego motion compensation Following the ego-motion compensation [4], we observe that static objects still have residual radial velocities when the vehicle rotate. Therefore we also account for the ego rotation. By doing so, we correctly remove both the translational and rotational ego-motion effects. Let $i$ denotes the $i$th radar point, $\bm{b}_{i}$ denotes the line-of-sight unit vector of radar point $p_{i}$. After ego motion compensation, the point velocity in the reference coordinate system is written as:
@@
Radar motion compensation After ego-motion compensation, dynamic objects can still appear displaced because radar detections are acquired at the sweep timestamp rather than at the keyframe timestamp. Unlike LiDAR methods that infer object motion via scene flow [44], radar Doppler directly measures velocity, capturing non-ego motion without learning. To mitigate the temporal offset, each radar point is translated according to its estimated motion over the time interval between the sweep timestamp $t_{s}$ and the keyframe timestamp $t_{k}$. The resulting motion-compensated point position is given by:
@@
Fig. 2 shows the effect of the proposed motion compensation in the bird’s-eye view (BEV). Because radar measures only radial velocity, tangential velocity is unobservable, causing residual distortion, especially for objects crossing the radar’s line of sight. This effect is strongest near the ego vehicle, so a nearby region is disabled (15 m ego-centric radius in this implementation).
@@ III-B Radar Informed Kalman filter
Motion Model We use a constant-velocity motion model. Let $k$ denotes the keyframe index, the state $\bm{x}_{k}=[x_{k},y_{k},v_{x,k},v_{y,k}]^{\top}$ denote the object position and planar velocity in the global coordinate system. $F$ denotes the state transition matrix, $\bm{Q}_{k}$ denotes the process noise. The predicted state is then given as:
@@
Radar Observation Model To handle noisy radar points, only radar points reflecting from the same object are used as observations. Let $p_{i}$ denote the $i$-th radar point, and let $\bm{p}^{(i)}_{s}\in\mathbb{R}^{3}$ denote its position, with measured radial velocity $v^{(i)}_{\mathrm{radar}}\in\mathbb{R}$, and let $v^{(i)}_{\mathrm{box}}\in\mathbb{R}$ denote the bounding box velocity projected onto the line-of-sight of point $i$. The associated radar point set is defined by two filters:
@@
To construct the radar observation matrix $\bm{H}_{r,k}$, where $r$ means observation from radar at the $k$-th frame, the line-of-sight unit vector of each associated radar point $i$ is defined from its position $\bm{p}_{s}^{(i)}\in\mathbb{R}^{3}$ in the radar sensor coordinate system $S_{i}$ as:
@@
where $\bm{b}_{i}\in\mathbb{R}^{3}$ is the line-of-sight direction in the radar coordinate system, $^{S_{i}}\bm{R}_{\mathrm{global},xy}\in\mathbb{R}^{2\times 3}$ extracts the planar rows of the global-to-radar rotation, and the position columns of $\bm{h}_{i}^{\top}$ are zero since only velocity enters the projection. Stacking $M$ such rows gives the joint observation matrix in (7).
@@
Radar-Informed Kalman Filter Update Given the radar observation, the tracking state is then updated using Kalman filter, as illustrated in Fig. 3, the radar-informed state is updated by:
@@ III-C Two-Stage Association
To reduce mismatches, which are the primary cause of identity switches, we propose a two-stage association that first performs a bidirectional cross-check association and then applies a radar association.
Cross-Check Association Instead of associating at a single timestamp (e.g., either comparing the forward prediction with the detection at time $t$, or comparing the track with the backward predicted detection at time t1t-1), we use both forward and backward predictions for mutual cross-checking and combine a speed similarity measure to further minimize the influence of object orientation. The cost function is defined as:
$cost=\alpha\frac{\|\bm{p}_{t}^{trk}-\bm{p}_{t}^{det}\|}{v_{t}^{trk}\Delta{t}+\epsilon}+\beta\frac{\|\bm{p}_{t-1}^{trk}-\bm{p}_{t-1}^{det}\|}{v_{t}^{det}\Delta{t}+\epsilon}+\delta\frac{|v^{trk}-v^{det}|}{\max(v^{trk},v^{det})+\epsilon}$ (10)
@@
Radar Association To recover the detector missed objects, unmatched tracked objects are associated with radar measurements in a second stage.
We first remove the points that close to detections to avoid duplicate associations, then select candidate radar points for unmatched tracked objects by three filters:
$\mathcal{R}^{t}_{\mathrm{assoc}}=\bigl\{\,p_{i}\in\mathcal{R}_{k}\;\big|\;p_{i}\in\beta\hat{\mathcal{B}}^{t}_{k}\;\wedge\;|v^{(i)}_{\mathrm{radar}}-v^{(i)}_{\mathrm{pred}}|\leq\delta_{v}\;\wedge\;\|\bm{u}^{(i)}_{\mathrm{radar}}\|>v_{\min}\,\bigr\}$ (11)
where
  • $\beta\hat{\mathcal{B}}^{t}_{k}$ restricts radar points to the track predicted bounding box, $\beta$ denotes the inflation rate.
  • $\delta_{v}$ removes outliers based on the radial velocity difference between each radar point $v^{(i)}_{\mathrm{radar}}$ and the predicted track velocity $v^{(i)}_{\mathrm{pred}}$ projected onto the radar point’s line-of-sight direction.
  • $v_{\min}$ filters out stationary clutter.
@@ IV EXPERIMENTS SETUP
Evaluation Metrics We follow the official nuScenes tracking tracking challenge [21] and report AMOTA. For all metrics, true positives are obtained by applying a 2 m threshold to the 2D center distance on the ground plane. AMOTA and MOTAR are defined as:
@@
Where $n$ denotes the number of recall sampling points, and $r$ is the recall threshold. $P$ refers to the number of ground-truth positives for the current class. ($\mathrm{IDS}$,($\mathrm{FP}$$\mathrm{FN}$) are the number of (identity switches, false positives, false negatives).
@@ V Results
We report our results on the TruckScenes validation set in Table I. The proposed RadarMOT achieves 33.3% in terms of overall AMOTA, an absolute improvement of 6.7% over the MCTrack baseline, and a relative 30% reduction in Identity Switches (IDS). We found that the distance-based association method is preferable in highly dynamic scenarios. We observed that CenterPoint achieves more true positives and fewer false negatives than MCTrack, but at the cost of a substantially increased number of IDS. Thanks to radar-refined state estimation, RadarMOT combines the strengths of both methods without introducing additional false positives. To the best of our knowledge, this is the first 3D MOT investigation on the TruckScenes, offering a conservative baseline for the community.
@@
Figure 4: Comparison across range on TruckScenes validation set.
Adverse Conditions We also evaluate the performance under adverse conditions in different weather, lighting, and driving areas, see Table III. RadarMOT achieves consistent absolute AMOTA improvements over the baseline MCTrack: +10.3% in fog, +10.8% at night, and +9.1% on highways. Contrary to our initial hypothesis, RadarMOT performs worse than the baseline in snow.
@@
(a) MCTrack
(b) RadarMOT (ours)
(c) MCTrack
(d) RadarMOT (ours)
Figure 5: Qualitative comparison on TruckScenes. Green bounding boxes indicate the ground truth, and same‑colored, numbered bounding boxes and trajectories represent individual tracking results. Images are displayed for context.
Qualitative Rerults Fig. 5 illustrates the robustness of our method in highway scenarios under challenging weather and lighting conditions, including nighttime driving, long-range perception in fog.

### 21 Apr 2026 v2

@@ Abstract
The challenge of 3D multi-object tracking is achieving robustness in real-world applications, for example under adverse conditions and maintaining consistency as distance increases. To overcome these challenges, sensor fusion approaches that combine LiDAR, cameras, and radar have emerged. However, existing multimodal methods usually treat radar as another learned feature inside the network. When the overall model degrades in difficult environments, the robustness advantages that radar could provide are also reduced. In this paper we propose RadarMOT, a radar-informed 3D multi-object tracking framework that explicitly uses radar point clouds as additional observations to refine state estimation and recover objects missed by the detector at long ranges. Evaluations on the MAN-TruckScenes dataset show that RadarMOT consistently improves the Average Multi-Object Tracking Accuracy (AMOTA) by 12.7% at long range and up to 10.3% in adverse weather. The code will be available at https://github.com/bingxue-xu/radarmot
@@ I Introduction
Safe 3D perception requires a system to continuously understand what is around it and where it is going before acting. This is the task of 3D multi-object tracking (3D MOT): maintaining identities and motion trajectories of detected objects across time. Trajectory prediction and motion planning depend directly on the quality of these estimates. 3D MOT operates by detecting objects independently in each frame and then associating them across time to form trajectories. This makes 3D MOT performance fundamentally dependent on two things: (i) how accurately motion is estimated (state estimation accuracy), and (ii) how reliably detections are matched with tracked objects (data association robustness).
@@
To reduce the computation burden on the GPU, we revisit both the radar sensing mechanism and Kalman filter-based 3D MOT frameworks [33, 31, 3, 24, 43]. We first incorporate radar point clouds as an additional measurement in the Kalman filter update, which consistently boosts Average Multi-Object Tracking Accuracy (AMOTA) and lowers false positives, indicating a more accurate state estimation. Next, we address motion blur during data preprocessing by leveraging radial velocity. Based on the observation that the distance-based tracker CenterPoint [37] yields more true positives than MCTrack [31], we further extend data association into a two-stage scheme that utilizes bidirectional distances with radar association, thereby stabilizing the tracking system.
Our proposed method RadarMOT achieves strong and consistent improvements on the MAN-TruckScenes dataset (TruckScenes), demonstrating that leveraging radar physical measurements in the tracking stage can substantially increase robustness when LiDAR and cameras degrade. Overall, RadarMOT improves AMOTA by +6.7% over the MCTrack baseline and reduces Identity Switches (IDS) by 30%, with the largest gains appearing at long range and under adverse conditions. The contributions of this paper are:
@@
  • We formulate a radar-informed Kalman filter update that uses associated radar radial velocities to refine tracked objects’ velocity and stabilize trajectories.
@@ II RELATED WORK
Tracking by Detection TBD divides the tracking task into four independently modules: detection, motion model, data association, and lifecycle management [24]. AB3DMOT [33] established a strong baseline by coupling a Kalman filter [25] with Intersection over Union (IoU)-based Hungarian matching [9]. The following methods have progressively refined the association strategy, from generalized IoU to normalization [47], lifting it to 3D [24], and combining distance with  IoU [31]. Another representative approach, CenterPoint [37], replaces  IoU with the Euclidean distance combined with greedy matching, achieving strong performance on nuScenes [2] and Waymo [27]. [3] replaces the Euclidean distance with the Mahalanobis distance. Other methods [13, 12, 43, 16] enhanced the motion model within the Kalman filter. SimpleTrack [24] proposes a two-stage association. MCTrack [31], consolidates these advances into a unified and general framework that refines the Kalman filter and incorporates both  IoU and distance into the data association stage, achieving state-of-the-art performance across the KITTI [6], nuScenes [2], and Waymo [27] datasets. We therefore adopt MCTrack as our baseline.
@@ III-A Motion Compensation
Ego motion compensation Following the ego-motion compensation [4], we observe that static objects still have residual radial velocities when the vehicle rotate. Therefore we also account for the ego rotation. By doing so, we correctly remove both the translational and rotational ego-motion effects. Let $i$ denote the $i$th radar point, $\bm{b}_{i}$ denotes the line-of-sight unit vector of radar point $p_{i}$. After ego motion compensation, the point velocity in the reference coordinate system is written as:
@@
Radar motion compensation After ego-motion compensation, dynamic points can still appear displaced because radar points are acquired at the sweep timestamp rather than at the keyframe timestamp. Unlike LiDAR methods that infer object motion via scene flow [44], radar Doppler directly measures velocity, capturing non-ego motion without learning. To mitigate the temporal offset, each radar point is translated according to its estimated motion over the time interval between the sweep timestamp $t_{s}$ and the keyframe timestamp $t_{k}$. The resulting motion-compensated point position is given by:
@@
Fig. 2 shows the effect of the proposed motion compensation in the bird’s-eye view (BEV). Because radar measures only radial velocity, tangential velocity is unobservable, causing residual distortion, especially for objects crossing the radar’s line of sight. This effect is strongest near the ego vehicle, so a nearby region is discarded (15 m ego-centric radius in this implementation).
@@ III-B Radar Informed Kalman filter
Motion Model We use a constant-velocity motion model, Let $k$ denote the keyframe index. The state $\mathbf{x}_{k}=[x_{k},y_{k},v_{x,k},v_{y,k}]^{\top}$ denotes the object position and planar velocity in the global coordinate system. $F$ denotes the state transition matrix, $\bm{Q}_{k}$ denotes the process noise. The predicted state is then given as:
@@
Radar Observation Model To handle noisy radar points, only radar points likely to be reflecting from the same object are used as observations. Let $p_{i}$ denote the $i$-th radar point, and let $\bm{p}^{(i)}_{s}\in\mathbb{R}^{3}$ denote its position, with measured radial velocity $v^{(i)}_{\mathrm{radar}}\in\mathbb{R}$, and let $v^{(i)}_{\mathrm{box}}\in\mathbb{R}$ denote the bounding box velocity projected onto the line-of-sight of point $i$. The associated radar point set is defined by two filters:
@@
To construct the radar observation matrix $\bm{H}_{r,k}$, where $r$ means observation from radar at the $k$-th frame, the line-of-sight unit vector $\bm{b}_{i}$ of each associated radar point $i$ is defined in the radar sensor coordinate system $S_{i}$ as:
@@
where $^{S_{i}}\bm{R}_{\mathrm{global},xy}\in\mathbb{R}^{2\times 3}$ extracts the planar rows of the global-to-radar rotation, and the position columns of $\bm{h}_{i}^{\top}$ are zero since only velocity enters the projection. Stacking $M$ such rows gives the joint observation matrix:
@@
Radar-Informed Kalman Filter Update Given the radar observation, the tracking state is then updated using a Kalman filter, as illustrated in Fig. 3, the radar-informed state is updated by:
@@ III-C Two-Stage Association
To reduce mismatches, which are the primary cause of identity switches in tracked objects, we propose a two-stage association that first performs a bidirectional cross-check association and then applies a radar association.
Cross-Check Association Instead of associating at a single timestamp (e.g., either comparing the forward prediction with the detection at time $t$, or comparing the track with the backward predicted detection at time t1t-1), we use both forward and backward predictions for mutual cross-checking and combine a speed similarity to further minimize the influence of object orientation. The cost function is defined as:
$cost=\frac{\alpha\|\bm{p}_{t}^{trk}-\bm{p}_{t}^{det}\|}{v_{t}^{trk}\Delta{t}}+\frac{\beta\|\bm{p}_{t-1}^{trk}-\bm{p}_{t-1}^{det}\|}{v_{t}^{det}\Delta{t}}+\frac{\delta|v^{trk}-v^{det}|}{\max(v^{trk},v^{det})}$ (10)
@@
Radar Association To recover the objects missed by the detector, unmatched tracked objects are associated with radar measurements in a second stage.
We first remove the points that are close to detections to avoid duplicate associations and remove stationary clutter, then select candidate radar points for unmatched tracked objects by position and velocity similarity.
$\mathcal{R}^{t}_{\mathrm{assoc}}=\bigl\{\,p_{i}\in\mathcal{R}_{k}\;\big|\;p_{i}\in\beta\hat{\mathcal{B}}^{t}_{k}\;\wedge\;|v^{(i)}_{\mathrm{radar}}-v^{(i)}_{\mathrm{pred}}|\leq\delta_{v}\,\bigr\}$ (11)
where $\beta\hat{\mathcal{B}}^{t}_{k}$ restricts radar points to the track predicted bounding box, $\beta$ denotes the inflation rate. $\delta_{v}$ removes outliers based on the radial velocity difference between each radar point $v^{(i)}_{\mathrm{radar}}$ and the predicted track velocity $v^{(i)}_{\mathrm{pred}}$ projected onto the radar point’s line-of-sight direction.
@@ IV EXPERIMENTS SETUP
Evaluation Metrics We follow the official nuScenes tracking challenge [21] and report AMOTA. For all metrics, true positives are obtained by applying a 2 m threshold to the 2D center distance on the ground plane. AMOTA and MOTAR are defined as:
@@
where $n$ denotes the number of recall sampling points, and $r$ is the recall threshold. $P$ refers to the number of ground-truth positives for the current class. ($\mathrm{IDS}$$\mathrm{FP}$$\mathrm{FN}$) are the number of (identity switches, false positives, false negatives).
@@ V Results
We report our results on the TruckScenes validation set in Table I. Our proposed method RadarMOT achieves 33.3% in terms of overall AMOTA, an absolute improvement of 6.7% over the MCTrack baseline, and a relative 30% reduction in IDS. We found that the distance-based association method is preferable in highly dynamic scenarios. We observed that CenterPoint achieves more true positives and fewer false negatives than MCTrack, but at the cost of a substantially increased number of IDS. Thanks to the bidirectional distance association and radar-refined state estimation, RadarMOT combines the strengths of both methods without introducing additional false positives. To the best of our knowledge, this is the first 3D MOT investigation on the TruckScenes, offering a conservative baseline for the community.
@@
Figure 4: Comparison across range and adverse conditions on TruckScenes validation set.
Adverse Conditions We also evaluate the performance under adverse conditions in different weather, lighting, and driving areas, see Table III and Fig. 4. RadarMOT achieves consistent absolute AMOTA improvements over the baseline MCTrack: +10.3% in fog, +10.8% at night, and +9.1% on highways. Contrary to our initial hypothesis, RadarMOT performs worse than the baseline in snow.
@@
(a) MCTrack
(b) RadarMOT (ours)
Figure 5: Qualitative comparison on TruckScenes in fog and at long distances. Green bounding boxes indicate the ground truth, and same‑colored, numbered bounding boxes and trajectories represent individual tracking results. Images are displayed for context.
Qualitative Rerults Fig. 5 illustrates the robustness of our method in highway scenarios under challenging foggy weather and lighting conditions.